Lisa Lasher buys 400 shares of stock on margin at $30 per share. If the margin requirement is 40 percent, how much must the stock rise for her to realize a 15-percent return on her invested funds? (Ignore dividends, commissions, and interest on borrowed funds.) Round your answer to the nearest cent.

$  
Answer #1

Margin requirement = 40%

Initial Investment = number of shares * price per share * margin requirement

Initial Investment = [$30 * 400 * 0.40] = $4,800

To realize a 15-percent return the value of position in stock must rise by ($4800*15%) = $720

The dollar must increase by ($720 / 400 share) = $1.80

Answer: Total Increase = $720

Per share increase = $1.80